# pH-Driven Fluorescence in Solid-State Lysine

### v1

@@ Abstract
There is currently a growing interest in understanding the origins of intrinsic fluorescence as a way to design non-invasive probes for biophysical processes. In this regard, understanding how pH influences fluorescence in non-aromatic biomolecular assemblies is key to controlling their optical properties in realistic cellular conditions. Here, we combine experiments and theory to investigate the pH-dependent emission of solid-state L-Lysine (Lys). Lys aggregates prepared at different pH values using HCl and H2SO4 exhibit protonation- and counterion-dependent morphology and fluorescence, as shown by microscopy and steady-state measurements. We find an enhancement in the fluorescence moving from acidic to basic conditions. To uncover the molecular origin of these trends, we performed non-adiabatic molecular dynamics simulations on three Lys crystal models representing distinct protonation states. Our simulations indicate that enhanced protonation under acidic conditions facilitates non-radiative decay via proton transfer, whereas basic conditions favor radiative decay. These findings highlight pH as a key factor tuning fluorescence in Lys assemblies, offering insights for designing pH responsive optical materials based on non-aromatic amino acids.
keywords:
@@
Fluorescence in biological materials is typically attributed to aromatic amino acids due to their conjugated -systems and favorable photophysical properties1. However, recent findings have revealed intrinsic photoluminescence in non-aromatic amino acids and peptides, particularly under aggregation in solution or in the solid state.2, 3, 4, 5, 6, 7, 8 A key mechanism proposed to underlie this phenomenon is aggregation-induced emission (AIE),9 wherein non-emissive species in dilute solution become strongly emissive upon aggregation. While AIE is well established in synthetic organic luminogens,10, 11, 12 its manifestation in non-aromatic biomolecular assemblies, such as peptide and amino acid aggregates, has opened new directions for exploring photophysics in biologically relevant systems.13, 14, 15, 16
@@
Hydrogen bonding plays a central role in mediating luminescence in non-aromatic systems, influencing both the extent of electron delocalization and which vibrational modes lead to non-radiative decay on the excited state17, 2, 18, 6. Environmental factors such as pH, solvent polarity, and ionic strength can profoundly affect these interactions, often altering morphology and photophysical properties19, 20. However, how these factors reshape the excited state (ES) potential energy landscape remains unclear21. While peptide self-assembly under neutral pH has been extensively studied,22, 23, 24 less attention has been given to extreme pH conditions, which can drastically alter conformation, charge distribution, and intermolecular interactions25, 21. In this regard, pH is known to play an important role in tuning fluorescence in aromatic or more generally conjugated molecular systems26, 27, 28, 29. Thus, the potential of designing novel probes based on intrinsic fluorescence as indicators of pH in optical sensing and imaging, could open up new applications in biological and material sciences29, 30.
@@
In this context, L-Lysine (Lys), a non-aromatic amino acid with a protonatable side chain, provides a compelling model to probe pH-dependent luminescence. Depending on pH, Lys adopts distinct protonation states (see Fig. 1),31 which in turn strongly influence its self-assembly and crystal packing, thereby modulating its photophysical properties.32, 4. Similar effects have been observed in poly-lysine systems, where pH influences aggregate morphology and optical response.33, 34, 35. Recent work by Stagi et al.32 combined experiments and theory to investigate the photophysical properties of self-assembled Lys in aqueous solution. Absorption spectra showed a distinct band at 278 nm, attributed to electronic transitions in hydrogen bonded Lys clusters. Classical molecular dynamics (MD) simulations further confirmed that Lys forms clusters of varied size through hydrogen bonds (HBs) and hydrophobic interactions, with the former playing a dominant role. Fluorescence measurements in solution exhibit strong pH dependence: emission is enhanced under neutral and basic conditions but quenched at low pH, where high protonation is proposed to inhibit functional group interactions, reducing emission efficiency.32 However, the precise ES relaxation pathways that affect fluorescence emission in titratable amino acids, such as lysine, remains poorly understood.
@@
In this contribution, by combining both experiments and theory, we address a critical gap in our understanding of how pH affects non-aromatic fluorescence using solid-state lysine as a model system. Solid-state Lys aggregates were prepared at four pH values (1, 4, 7, and 10) to probe how the protonation influences both morphological and photoluminescence properties. Since pH was adjusted with either a monoprotic (HCl) or a diprotic (H2SO4) acid, we also assessed the impact of the counterion on these properties. Morphology was characterized by optical and scanning electron microscopy (SEM), while photoluminescence was studied via confocal microscopy, steady-state fluorescence, and external quantum efficiency (EQE) measurements. Our results show that both protonation state and the selected acid significantly influence aggregate morphology and fluorescence, establishing these as key parameters for tuning optical properties. To uncover the molecular origins behind these trends, we performed non-adiabatic molecular dynamics (NAMD) simulations to follow the ES dynamics of Lys. While direct simulation of the experimental aggregates is precluded by the lack of precise structural data, we can instead model crystal structures of Lys representing distinct protonation states (Lys2+, Lys+, and Lys0 in Fig. 1).36, 37, 38 These models offer well-defined frameworks to interpret fluorescence behavior and examine ES processes. Our simulations indicate that fluorescence decreases from low to high protonation states (), consistent with our experimental data and prior reports32, 4. Mechanistically, increased protonation promotes proton transfer (PT) processes that enable non-radiative decay, causing fluorescence quenching under acidic conditions.
@@
As mentioned above, we investigated the morphology of solid-state Lys aggregates formed at pH 1, 4, 7, and 10, adjusted using either HCl or H2SO4, by optical and scanning electron microscopy. The characterization was performed on Lys samples prepared in water at the concentration of g/mL ( mol/L) corresponding to the maximum of its solubility. This concentration–which is approximately tenfold higher than the critical aggregation concentration (CAC)–ensures the formation of aggregated clusters (see Fig. S1 and discussion in the Supporting Information (SI)).
@@
Fig. 2 shows clear morphological differences depending on both acid type and pH. Under strongly acidic conditions (pH = 1), both HCl- and H2SO4-treated samples initially formed oily films at room temperature, with no macroscopically ordered structures observed over time. However, upon overnight cooling to , well-ordered crystals emerged (Fig. 2a, Lys + HCl; Fig. 2e, Lys + H2SO4). Optical microscopy revealed distinct morphologies: HCl-treated samples formed smaller, densely packed crystals, likely due to enhanced nucleation promoted by chloride ions. In contrast, H2SO4 promoted the formation of larger, elongated crystals, suggesting that sulfate anions favor fewer nucleation events, leading to more sparsely distributed structures. Reheating caused crystals to revert to oily films, preventing further morphological study.
@@
Overall, optical and scanning electron microscopy reveal that both pH and acid type influence Lys aggregate morphology. To investigate how these structural differences impact optical properties, we turned to fluorescence-based techniques, starting with confocal and fluorescence microscopy, followed by steady-state spectroscopy and quantum efficiency measurements to better quantify the pH dependent fluorescence emission.
@@
To quantify this, we collected steady-state fluorescence spectra across excitation wavelengths from 330-410 nm (10 nm step). The resulting spectra for Lys acidified with HCl and H2SO4 are shown in panels a-d and panels e-h of Fig. S4, respectively. Interestingly, we observed a linear correlation between excitation and emission wavelengths (Fig. S5 of the SI), consistent with previous findings in Lys aggregates in solution32, and poly-lysine polymers33, where this dependency was attributed to the ability of Lys to form aggregates of various size. Excitation-dependent fluorescence, or red-edge excitation shifts, has been observed in a variety of systems, including carbon nanodots39, sodium-uracil thin films40, self-assembled peptide nucleic acids41, and organic fluorescent nanoparticles42. Proposed mechanisms include heterogeneous aggregate populations39, slow environment relaxation around the excited fluorophore41, 40, and collective proton motions within HB networks that modulate the ES manifold42. These findings suggest that the excitation-dependent emission in our Lys samples likely arises from similarly heterogeneous supramolecular environments, shaped by protonation state and ionic interactions.
@@
To isolate the role of protonation and acid identity, we compared normalized fluorescence spectra at nm for all samples, as shown in Fig. 3e-f. Both acid treatments yielded broad emissions (350–550 nm), with emission maxima () centered between 400 nm and 420 nm, depending on pH and acid. For HCl (Fig. 3e), blue-shifts from 422 nm (pH 1, red line) to 403 nm (pH 10, cyan line), while for H2SO4 (Fig. 3f), the pH 1 sample shows the most blue-shifted emission ( nm), and all other pH values exhibit nearly identical spectra ( 414 nm). These subtle spectral differences align with SEM observations: H2SO4 yields more consistent fibrous morphologies across pH, while HCl shows greater variation, potentially explaining the progressive red-shift.
@@
While these spectra offer an initial view of how pH influences fluorescence, their normalization masks variations in absolute emission intensity. For this reason, we measured external quantum efficiency (Fig. 3g), since it accounts for light extraction efficiency43 and is therefore particularly relevant for assessing emission in solid-state devices44, 45, 46. Here, a striking pH dependence emerges: EQE increases from at pH 1 for both acid treatments and rises to nearly 5 at pH 7 for Lys with HCl (green plot) and at pH 7-10 for Lys with H2SO4 (purple plot), reflecting higher emission efficiency under low protonation. These trends are consistent with previous studies on solution-phase Lys aggregates, which linked protonation to fluorescence quenching32. Notably, the two acids impart distinct efficiencies: with H2SO4, EQE increases from pH 1 to pH 7 and then plateaus at pH 10, whereas with HCl the efficiency peaks at pH 7 and declines at pH 10. This divergence underscores the additional role of the counterion, beyond protonation alone, in dictating aggregate structure and, consequently, solid‑state fluorescence efficiency.

### v2

# Investigating the Role of pH and Counterions in the Intrinsic Fluorescence of Solid-State L-Lysine

@@ Abstract
There is currently a growing interest in understanding the origins of intrinsic fluorescence as a way to design non-invasive probes for biophysical processes. In this regard, understanding how pH influences fluorescence in non-aromatic biomolecular assemblies is key to controlling their optical properties in realistic cellular conditions. Here, we combine experiments and theory to investigate the pH-dependent emission of solid-state L-Lysine (Lys). Lys aggregates prepared at different pH values using HCl and H2SO4 exhibit protonation- and counterion-dependent morphology and fluorescence, as shown by microscopy and steady-state measurements. We find an enhancement in the fluorescence moving from acidic to basic conditions. To uncover the molecular origin of these trends, we performed non-adiabatic molecular dynamics simulations on three Lys crystal models representing distinct protonation states. Our simulations indicate that enhanced protonation under acidic conditions facilitates non-radiative decay via proton transfer, whereas basic conditions favor radiative decay. Our combined experimental-theoretical work highlights pH and counterion identity as key factors tuning fluorescence in Lys assemblies, offering insights for designing pH responsive optical materials based on non-aromatic amino acids.
@@ 1 Introduction
Fluorescence in biological materials is typically attributed to aromatic amino acids due to their conjugated -systems and favorable photophysical propertiesLakowicz_2006. However, recent studies have revealed that non-aromatic amino acids and peptides can also exhibit intrinsic photoluminescence, particularly under aggregation in solution or in the solid state.Pinotsi2016, Grisanti2020, arnon2021off, new_tang2021nonconventional, miron2023, Banerjee_2025, zhang2025_gluser A key mechanism proposed to underlie this phenomenon is aggregation-induced emission (AIE),Luo_2001 wherein non-emissive species in dilute solution become strongly emissive upon aggregation. While AIE is well established in synthetic organic luminogens,He_2018, Zhao_2020, yu_2020 its manifestation in non-aromatic biomolecular assemblies, such as peptide and amino acid aggregates, has opened new directions for exploring photophysics in biologically relevant systems.delmercato_2007, new_chen2018prevalent, Lapshina_2019, Zhang_2019
@@
Hydrogen bonding plays a central role in mediating luminescence in these non-aromatic systems, influencing both the extent of electron delocalization and the vibrational modes that promote non-radiative decay from the excited state (ES)SHUKLA_2004, Pinotsi2016, Stephens2021, miron2023. Environmental factors such as pH, solvent polarity, and ionic strength can profoundly affect these interactions, often altering morphology and photophysical propertiesFung_2006, Clementi_2019. However, how these factors reshape the ES potential energy landscape remains unclearWang_2025. While peptide self-assembly under neutral pH has been extensively studied,Wang_2018, Sarkhel_2020, McCahill_2024 less attention has been given to extreme pH conditions, which can drastically alter conformation, charge distribution, and intermolecular interactionsLi_2022, Wang_2025. In this regard, pH is known to play an important role in tuning fluorescence in aromatic or more generally conjugated molecular systemswhite1959effect, STEINER1963341, berezin2009ph, Han_2010. Thus, the potential of designing novel probes based on intrinsic fluorescence as indicators of pH in optical sensing and imaging, could open up new applications in biological and material sciencesHan_2010, Steinegger_2020.
@@
In this context, L-Lysine (Lys), a non-aromatic amino acid with a protonatable side chain, provides a compelling model to probe pH-dependent luminescence. Depending on pH, Lys adopts distinct protonation states (see Fig. 1),Nolting_2007 which in turn strongly influence its self-assembly and crystal packing, thereby modulating its photophysical properties.stagi2022, arnon2021off. Similar effects have been observed in poly-lysine systems, where pH influences aggregate morphology and optical response.Stagi_2021, stagi_2022_polyLys, Kastinen_2023_polyLys. Recent work by Stagi et al.stagi2022 combined experiments and theory to investigate the photophysical properties of self-assembled Lys in aqueous solution. Absorption spectra showed a distinct band at 278 nm, attributed to electronic transitions in hydrogen bonded Lys clusters. Classical molecular dynamics (MD) simulations further confirmed that Lys forms clusters of varied size through hydrogen bonds (HBs) and hydrophobic interactions, with the former playing a dominant role. Fluorescence measurements in solution exhibit strong pH dependence: emission is enhanced under neutral and basic conditions but quenched at low pH, where high protonation is proposed to inhibit functional group interactions, reducing emission efficiency.stagi2022 However, the precise ES relaxation pathways that affect fluorescence emission in titratable amino acids, such as lysine, remains poorly understood.
@@
In this work, by combining both experiments and theory, we address a gap in our understanding of how pH affects non-aromatic fluorescence using solid-state lysine as a model system. Solid-state Lys aggregates were prepared at four pH values (1, 4, 7, and 10) to probe how the protonation influences both morphological and photoluminescence properties. Since pH was adjusted with either a monoprotic (HCl) or a diprotic (H2SO4) acid, we also assessed the impact of the counterion on these properties. Morphology was characterized by optical and scanning electron microscopy (SEM), while photoluminescence was studied via confocal microscopy, steady-state fluorescence, and external quantum efficiency (EQE) measurements. Our results show that both protonation state, the choice of the acid and subsequently the counterion in the corresponding crystal, significantly influence aggregate morphology and fluorescence, establishing these as key parameters for tuning optical properties. To uncover the molecular origins behind these trends, we performed non-adiabatic molecular dynamics (NAMD) simulations to follow the ES dynamics of Lys. While direct simulation of the experimental aggregates is precluded by the lack of precise structural data, we can instead model crystal structures of Lys representing distinct protonation states (Lys2+, Lys+, and Lys0 in Fig. 1).Capasso_1983, wright_1962, Williams_2015 These models offer well-defined frameworks to interpret fluorescence behavior and examine ES processes. Our simulations indicate that fluorescence decreases from low to high protonation states (), consistent with our experimental data and prior reportsstagi2022, arnon2021off. In short, increased protonation promotes proton transfer (PT) processes that enable non-radiative decay, causing fluorescence quenching under acidic conditions.
2 Methods
2.1 Experimental Details
2.1.1 Chemicals and Samples Preparation
L-Lysine (Lys) monohydrate, HCl (wt) and H2SO4 (wt) were purchased by Thermo Fisher Scientific (Milan, Italy). All solutions were prepared by weight dissolving Lys in doubly distilled water; their pH was adjusted using HCl (wt) and H2SO4 (wt). pH was measured by pH-meter (Mettler Toledo FiveEasy Plus pH meter FP20, equipped with a micro electrode). Experimental conditions were verified to be above the critical aggregation concentration (CAC) of Lys, determined through fluorescence probe assays, as reported in details in the Supporting Information (SI).
2.1.2 Scanning Electron Microscopy
Morphological characterization of the samples was performed using field-emission scanning electron microscopy (ThermoFisher Scientific Phenom Pro G6). Aliquots of 10 L Lys solutions were drop-cast onto aluminum stubs and air-dried at room temperature. Samples were sputter-coated with a 7 nm gold layer (25 mA, 75 s) prior to imaging. Images were acquired in high vacuum mode at an accelerating voltage of 15 kV using the Secondary Electron Detector.
2.1.3 Optical Microscopy
Samples were prepared by drop-casting 20 L of each Lys solution onto glass microscope slides, which were then air-dried overnight at room temperature. The dried samples were examined using an Optech BM80Pol microscope (Milan, Italy) under both cross-polarized and bright-field light.
2.1.4 Confocal Microscopy
Samples of Lys solutions at various protonation states were drop-cast onto microscope slides and allowed to air-dry overnight. The dried samples were analyzed on the MICA microhub (Leica, Wetzlar, Germany) platform using and magnifications. Fluorescence images were acquired using filters in the spectral regions of blue (, ), green (, ), and red (, ).
2.1.5 Fluorescence Spectroscopy
Fluorescence spectra were acquired at room temperature on a Jasco spectrofluorophotometer (Model FP-8550), equipped with a continuous output Xenon arc lamp. Measurements in solution were conducted in a quartz cell with a 1.0 cm path length. Solid state fluorescence and photoluminescence quantum yields (PLQYs) were evaluated using an ILF-835 integrating sphere. The absolute method was employed for PLQY determination thereby overcoming the issue of unavoidable differences in the relative emission intensities arising from poorly reproducible drop-casted amino acid clusters. Fluorescence spectra of solid-state Lys were collected up to 600 nm by scanning the 330–410 nm excitation range in 10 nm increments. All acquisitions were averaged over five measurements after spectral correction.
2.2 Computational Details
2.2.1 Classical Molecular Dynamics
Lys systems in different protonation states were modeled using crystallographic structures of L-Lysine sulfate (LLS), L-Lysine monohydrochloride (LLMHCl), and L-Lysine hemihydrate (LLH). Supercells of appropriate size were built for each system: 1588 for LLS, 11511 for LLMHCl, and 7116 for LLH. Classical molecular dynamics (MD) simulations were performed using the AMBER MD package version 2023amber2005, ambertools2023 with the Generalized Amber Force Field (GAFF) gaff_wang2004 applied to all molecular mechanics (MM) atoms. For each supercell, we first performed energy minimization to remove any unfavorable contacts. Minimized systems were then equilibrated for 10 ns in the NVT ensemble at 300 K (2 fs time step, Langevin thermostatIzaguirre_2001 with ps-1). Root mean square deviation (RMSD) relative to the first frame and potential energy were monitored along the simulations, as shown in Fig. S1 of the SI, confirming that the systems equilibrate within 10 ns. For each system, 50 independent frames were extracted from the corresponding MD production and used as starting geometries for the subsequent quantum mechanics/molecular mechanics (QM/MM) simulationsqmmm_warshel_levitt_1976, qmmm_brunk2015mixed.
2.2.2 Ground State QM/MM
For each crystal, we selected a suitable QM region (see Fig. 5a–c) which was then equilibrated at 300 K using ORCA 5.0.4orca_neese2012, orca_neese2018, orca_neese2020, orca_neese2022, orca_neese2023shark (QM) interfaced with AMBER 2023amber2005, ambertools2023 (MM). The QM region was treated using the CAM-B3LYP exchange-correlation functionalcamb3lyp_yanai2004 and the def2-SVP basis setdef2_weigend2005, def2_hellweg2015 with the RIJ-COSX approximationrijcosx_neese2009 to accelerate Coulomb and exact exchange integral evaluations. A first equilibration of 1 ps was performed in the NVT ensemble with tight thermostat coupling ( ps-1, 0.5 fs step, no bond constraints), followed by 1 ps with a standard thermostat coupling ( ps-1). Final nuclear coordinates and velocities were used as initial conditions for the non-adiabatic molecular dynamics (NAMD) simulations described below.
The thermally equilibrated structures were further used to compute vertical excitation energies at the time-dependent density functional theory (TDDFT) level. These calculations were performed using the same ORCA/AMBER QM/MM framework described above, but with no MD propagation (0 MD steps). For each Lys crystal system, the ten lowest singlet excited states were evaluated to extract the absorption spectra of the different crystal structures. Molecular orbitals were computed using a post-processing tool available in ORCA.
2.2.3 Non-adiabatic Molecular Dynamics
NAMD simulations were performed using the Trajectory Surface Hopping (TSH) method within the SHARC formalismsharc_richter2011, sharc_mai2018, which implements Tully’s Fewest Switches Surface Hopping (FSSH)tsh_tully1990molecular and allows the inclusion of time-dependent non-adiabatic couplings (TD-NACs). In this approach, nuclei evolve classically on electronic potential energy surfaces computed on-the-fly. For a comprehensive overview of TSH methodology, the reader is referred to Refs. tsh_barbatti2011nonadiabatic, malhado2014non, persico2014overview, tsh_crespo2018recent.
For each system, 50 trajectories were propagated including the first five singlet states () to adequately describe the lowest excited states relevant for fluorescence studies. All trajectories were initialized on the first excited state () and propagated in the NVE ensemble for 500 fs with 0.5 fs time step using the velocity Verlet integration algorithmverlet1967computer, swope1982computer. The ground-state QM/MM functional and basis set were retained. Excitation energies and oscillator strengths were calculated at the TDDFTtddft_casida1995 level within the Tamm-Dancoff approximation (TDA)tda_hirata1999, which improves TDDFT behavior near conical intersectionsCordova_2007, Tapavicza_2008, Peach_2012. NACs were computed via the Hammes-Schiffer–Tully schemetsh_hammes1994proton based on wavefunction overlap integrals between successive time steps. Electronic dynamics were integrated using 25 sub-steps per nuclear time step, and decoherence effects were included using the energy-based correction of Granucci et al.tsh_granucci2007critical with a 0.1 Hartree parameter. Surface hops to S0 were allowed when the energy gap between the current state and S0 was below 0.2 eV, consistent with prior TDDFT-based NAMD studies.plasser2014surface, wen2023excited, ibele2020excimer Kinetic energy was rescaled along the direction of the NAC vector to conserve energy during allowed hops; frustrated hops were discarded. Simulations were performed using SHARC 3.0.1SHARC3.0 for surface hopping, ORCA 5.0.4orca_neese2012, orca_neese2018, orca_neese2020, orca_neese2022, orca_neese2023shark for quantum chemical calculations, and TINKER 6.3.3tinker_ponder2004 for MM dynamics.
3 Results and Discussion
3.1 Experimental Results
3.1.1 Structural Characterization of Lysine Aggregates
@@
We begin by examining the morphology of solid-state Lys aggregates prepared at pH 1, 4, 7, and 10, with pH adjusted using either HCl or H2SO4, by optical and scanning electron microscopy. The characterization was performed on Lys samples prepared in water at the concentration of g/mL ( mol/L) corresponding to the maximum of its solubility. At this concentration, the experimentally measured pH is 11.8, approximately two units above the isoelectric point (9.74.) Under these conditions, the concentration is approximately tenfold higher than the CAC, ensuring the formation of aggregated clusters (see Fig. S2 and discussion in the SI).
@@
Fig. 2 shows clear morphological differences depending on both acid type and pH. Under strongly acidic conditions (pH = 1), both HCl- and H2SO4-treated samples initially formed oily films at room temperature, with no macroscopically ordered structures observed over time. This behavior can be rationalized by considering that, at this pH, Lys is in its double protonated state, which introduces strong electrostatic repulsion between molecules. The presence of counterions (Cl- or SO) can screen these interactions, enabling molecular association. However, at room temperature, the combination of high solubility,Amend_1997 thermal motion, and residual repulsion may hinder the establishment of long-range order. Upon overnight cooling to , reduced molecular motion enhances the stability of weak directional interactions, leading to the formation of well-ordered crystals (Fig. 2a, Lys + HCl; Fig. 2e, Lys + H2SO4), in line with previous studies on cooling-induced crystallizationChristopher_1998. Optical microscopy revealed distinct morphologies: HCl-treated samples formed thinner, densely packed crystals, whereas H2SO4 promoted larger, plate-like ones, suggesting that the two counterions influence nucleation behavior differently. Reheating caused crystals to revert to oily films, precluding further morphological study, highlighting the role of temperature in tuning the equilibrium between disordered and crystalline phases.Chen_2022, Monti_2023.
@@
3.1.2 pH-Dependent Fluorescence in Lysine Aggregates
Since both pH and acid type influence Lys aggregate morphology, and aggregate structure is known to affect the emergence and intensity of intrinsic fluorescence in non-aromatic systems,berger_2015, Pinotsi2016, stagi2022 we next investigated how these morphological variations affect the optical response. To establish this, we turned to fluorescence-based techniques, starting with confocal and fluorescence microscopy, followed by steady-state spectroscopy and quantum efficiency measurements to quantify the pH-dependent emission.
@@
To quantify this, we collected steady-state fluorescence spectra across excitation wavelengths from 330-410 nm (10 nm step). The resulting spectra for Lys acidified with HCl and H2SO4 are shown in panels a-d and panels e-h of Fig. S5, respectively. A linear correlation between excitation and emission wavelengths was observed (Fig. S6), consistent with previous reports on Lys aggregates in solutionstagi2022, and poly-lysine polymersStagi_2021, where this behavior was attributed to the ability of Lys to form aggregates of different size. Such excitation-dependent fluorescence, or red-edge excitation shift, has been observed in a variety of systems, including carbon nanodotsSharma_2016, sodium-uracil thin filmspale_2017, self-assembled peptide nucleic acidsberger_2015, and organic fluorescent nanoparticleslee2013. Proposed origins include heterogeneous aggregate populationsSharma_2016, slow relaxation of the local environment around the excited fluorophoreberger_2015, pale_2017, and collective proton motions within HB networks that modulate the ES manifoldlee2013. We propose that, in our Lys samples, the excitation-dependent emission likely arises from similarly heterogeneous supramolecular environments shaped by protonation state and ionic interactions.
@@
To isolate the role of protonation and acid identity, we compared normalized fluorescence spectra at nm for all samples, as shown in Fig. 4a-b. Both acid treatments yielded broad emissions (350–550 nm), with emission maxima () centered between 400 nm and 420 nm, depending on pH and acid. For HCl (Fig. 4a), blue-shifts from 422 nm (pH 1, red line) to 403 nm (pH 10, cyan line), while for H2SO4 (Fig. 4b), the pH 1 sample shows the most blue-shifted emission ( nm), and all other pH values exhibit nearly identical spectra ( 414 nm). These subtle spectral differences qualitatively correlate with the SEM observations. H2SO4-treated samples exhibit more homogeneous fibrous morphologies across pH, consistent with a uniform molecular organization and emission response. In contrast, the greater morphological variability observed for HCl-treated samples suggests a broader range of local packing arrangements, which may contribute to the observed spectral differences. Although this remains an empirical correlation, the SEM data provide qualitative insight into the molecular arrangement influencing fluorescence behavior, in line with previous reports correlating aggregate morphology and emission propertiesHong_2011.
@@
While these spectra offer an initial view of how pH influences fluorescence, their normalization masks variations in absolute emission intensity. To address this, we measured external quantum efficiency (Fig. 4c), which accounts for light extraction efficiencyPark_2024 and is therefore particularly relevant for assessing emission in solid-state devicesliu_2017, song_2018, Jang_2020. A clear pH dependence emerges: EQE increases from at pH 1 for both acid treatments to nearly 5 at pH 7 for Lys with HCl (green plot) and at pH 7-10 for Lys with H2SO4 (purple plot), reflecting higher emission efficiency under low protonation. These trends are consistent with previous studies on solution-phase Lys aggregates, which linked protonation to fluorescence quenchingstagi2022. Notably, the two acids impart distinct efficiencies: with H2SO4, EQE increases from pH 1 to pH 7 and then plateaus at pH 10, whereas with HCl the efficiency peaks at pH 7 and declines at pH 10. This divergence underscores the additional role of the counterion, beyond protonation alone, in dictating aggregate structure and, consequently, solid‑state fluorescence efficiency.
3.2 Theoretical Results
3.2.1 Excited State Simulations of Lysine Crystals